OpenStudy (calculusxy):

A cart is traveling at a constant velocity of 0.70 m/s forwards. A fan that can increase the car's velocity at a rate of 2.8 m/s^2 is suddenly turned on and accelerates the cart forwards for 3.5 seconds. How fast will the cart be moving after this time?

OpenStudy (calculusxy):

@SolomonZelman @mathstudent55

OpenStudy (shawn):

what formula do you know?

OpenStudy (calculusxy):

acceleration = change in velocity/change in time

OpenStudy (shawn):

that's the one

OpenStudy (shawn):

and Δvelocity = final velocity - initial velocity. You have everything. Just plug and chug to find final velocity

OpenStudy (calculusxy):

so these are the info that i have gathered so far: change in time = 3.5 seconds initial velocity = 0.70 m/s final velocity = ? acceleration = 2.8 m/s^2

OpenStudy (shawn):

yes

OpenStudy (calculusxy):

2.8 m/s^2 = (final velocity - 0.70 m/s) / 3.5 s

OpenStudy (shawn):

yes

OpenStudy (shawn):

now just do the algebra.

OpenStudy (calculusxy):

9.8 m/s = final velocity - 0.70 m/s 11.5 m/s = final velocity (but the answer is 11 m/s)

OpenStudy (calculusxy):

@Nnesha

OpenStudy (calculusxy):

nvm i got 10.5 and so it rounds to 11 m/s because of sigfigs

OpenStudy (calculusxy):

thanks for the help!

OpenStudy (moldybubblegum11):

\[\Huge\color{yellow}\star \]

OpenStudy (moldybubblegum11):

\[\Huge\color{orange}\checkmark \]